OpenStudy (anonymous):

Carolyn drove 150 miles in 3 hours.    Which rate represents this situation?

OpenStudy (texaschic101):

150 /3 = 50 /1 (150 miles in 3 hrs = 50 miles per her)

OpenStudy (texaschic101):

oops....typo...per HOUR

OpenStudy (anonymous):

Tiffany sold 4 paintings in 3.5 months. Which rate represents this situation

OpenStudy (texaschic101):

4/3.5 = 1.14/1 --- (4 paintings in 3.5 months = 1.14 paintings in 1 month)